a) Find 𝑑𝑦


𝑑π‘₯


𝑔𝑖𝑣𝑒𝑛 π‘‘β„Žπ‘Žπ‘‘ 𝑦 = (π‘₯


2 – 3x + 1)4


b) Differentiate 2


π‘₯(x)βˆ’3π‘₯+1



c) Intergrate 2x2 2

1
Expert's answer
2022-03-07T13:26:01-0500

a) "y=(x^2-3x+1)4"

we start by simplifying the equation

"y=4x^2-12x+4"

we apply the sum/difference rule

"dy\/dx =" "d\/dx(4x^2)-d\/dx(12x)+d\/dx(4)"

"=8x-12"

b. "y= 2x(x) -3(x)+1"

we start by simplifying the equation

"y=2x^2-3x+1"

to find the derivative we use sum difference rule

"dy\/dx= d\/dx(2x^2)-d\/dx(3x)+d\/dx(1)"

"=4x-3"

c. Incase the given equation is as follows

"\int( 2x^2+2)dx"

we apply the sum rule of integration

"=\int2x^2dx+\int2dx"

"=2x^3\/3+2x+C"

where C is the constant of integration.

Incase the given equation is as follows

= "\int (2x^2-2)dx"

we apply the difference rule of integration

"=\int2x^2dx-\int 2dx"

where C is the constant of integration.